Underwriting and Flotation Expenses

The Beranek Company, whose stock price is now $30, needs to raise $30 million in common stock. Underwriters have informed the firm's management that they must price the new issue to the public at $26 per share because of signaling effects. The underwriters' compensation will be 4% of the issue price, so Beranek will net $24.96 per share. The firm will also incur expenses in the amount of $170,000. How many shares must the firm sell to net $30 million after underwriting and flotation expenses? Do not round intermediate calculations. Round your answer to the nearest whole number.
Answer #1

Let the number of shares to be sold be x
Hence,
30*10^6=x*24.96-170000
=>x=(30*10^6+170000)/24.96
=>x=1208733.9743590

Answer is 1208734